# Exploring entanglement in finite-size quantum systems with degenerate ground state

### v1

@@ Results
where the index $d$ denotes degenerate eigenstates, $\mu$ stands for projection of the spin operator and $i$ is the site index. For the models we consider the local magnetization can be calculated exactly. In Fig.8 we present these exact values of the local magnetization with lines. Blue color denotes the range of the magnetic fields with degenerate solutions. These results fully agree with that known from the literature Pfeuty .
@@
Conclusion
@@
In this work we address the problem of characterizing the quantum systems in degenerate ground states with a special focus on entanglement entropy properties. On the basis of the linearly independent degenerate eigenstates calculated with exact diagonalization we propose to generate a finite ensemble of their linear combinations with Haar measure. This allows to probe the entanglement distributions of the ground wave functions. The primers we consider in this work show that such distributions are problem specific. For instance, in the case of the transverse-field Ising model the calculated values of the von Neumann entropy are non-uniformly distributed in the wide range between 0 and 1. At the same time, for the quantum system with non-collinear order the entanglement distribution reveals a more localized Gaussian-like shape. The found enhancement of the mean entropy of the degenerate eigenstate ensemble in the spin spiral phase can be explained by the changes in three- and two-spin correlations which demonstrate different trends at increasing the external magnetic field. To establish a connection between our theoretical results and real experiments we analyze the important details concerning estimation of the observables in the case of the degenerate quantum systems.
Acknowledgements
@@
The research funding from the Ministry of Science and Higher Education of the Russian Federation (Ural Federal University Program of Development within the Priority-2030 Program, project 4.72) is gratefully acknowledged. Quantum simulations were performed on the Uran supercomputer at the IMM UB RAS.

### v2

@@ V Single-shot description of the degenerate systems
where the index $d$ denotes degenerate eigenstates, $\mu$ stands for projection of the spin operator and $i$ is the site index. For the Ising and DMI models we consider the local magnetization can be calculated exactly. In Fig.10 we present these exact values of the local magnetization with lines. Blue color denotes the range of the magnetic fields with degenerate solutions. These results fully agree with that known from the literature Pfeuty ; qskyrmion .
@@
VI Simple fermionic and bosonic systems with degeneracy of the ground state
To complete our study of the quantum systems with degenerate ground state in this section we extend our consideration onto systems interacting fermions or bosons that attract considerable attention in modern experiments on ultra-cold atoms Ultracold . As we will demonstrate the simplest two-site Fermi-Hubbard and Bose-Hubbard models can reveal interesting entanglement properties in some limiting cases.
First we discuss the results obtained for two-site fermionic Hubbard model Hubbard that is given by
$H_{\rm FH}=-t\sum_{ij\sigma}c^{\dagger}_{i\sigma}c_{j\sigma}+U\sum_{i}n_{i\uparrow}n_{i\downarrow}-\mu\sum_{i\sigma}n_{i\sigma},$ (15)
where $t_{ij}$ is the hopping integral, $U$ stands for the on-site Coulomb interaction, $\mu$ is chemical potential. $c^{\dagger}_{i\sigma}$ ($c_{i\sigma}$) and $n_{i\sigma}$ are the fermionic creation (annihilation) and particle number operators, respectively. Each site can be in four different states: empty |$\ket{0}$, occupied by an electron with spin up $\ket{\uparrow}$, occupied by an electron with spin down $\ket{\downarrow}$ and occupied by two electrons with opposite spins $\ket{\uparrow\downarrow}$.
Among possible solutions of the Hubbard model we are primarily interested in exploring those with the degeneracy of the ground state. For instance, this is the case when there is single electron in two-site system, whose ground state is described by the following eigenstates
|$\ket{\psi_{\uparrow}}=\frac{1}{\sqrt{2}}(\ket{\uparrow 0}+\ket{0\uparrow}),$ (16)
|$\ket{\psi_{\downarrow}}=\frac{1}{\sqrt{2}}(\ket{\downarrow 0}+\ket{0\downarrow}).$ (17)
The corresponding ground state energy is equal to $-t-\mu$. For single-site subsystem the entanglement entropy of each state is equal to 1 and relates to our uncertainty to find the electron with spin up or spin down at the specific site. In turn, the von Neumann entropy calculated for an arbitrary superposition of these degenerate wave functions, $\alpha_{0}\ket{\psi_{\uparrow}}+\alpha_{1}\ket{\psi_{\downarrow}}$ is likewise equal to 1 independent on the choice of the coefficients. Thus, the degeneracy relating to the spin channels doesn’t affect quantum correlation between different sites in the considered single-particle case.
Next we are going to discuss the two-site system of interacting bosons that are described with the following Hubbard model
$H_{\rm BH}=-t\sum_{ij}a^{\dagger}_{i}a_{j}+\frac{U}{2}\sum_{i}n_{i}(n_{i}-1)-\mu\sum_{i}n_{i},$ (18)
where $a^{\dagger}_{i}$ ($a_{i}$) and $n_{i}$ are the bosonic creation (annihilation) and particle number operators, respectively.
[…]
For $N$=1 there is two-fold degenerate ground state with $E_{0}=E_{1}=-\mu$ and the corresponding wave functions are |$\ket{\psi_{0}}=\frac{1}{\sqrt{2}}(\ket{01}+\ket{10})$ and |$\ket{\psi_{1}}=\frac{1}{\sqrt{2}}(\ket{01}-\ket{10})$. While each of these wave functions features 50% probability to find the boson at the specific site and the maximal entanglement entropy value of 1, this is not the case for their superpositions within the degeneracy manifold. For instance, the combination $\frac{1}{\sqrt{2}}(\ket{\psi_{0}}+\ket{\psi_{1}})$ gives zero entanglement, since the boson occupies the first site. The distribution of the entanglement is almost the same as found in the case of the Ising model Fig.4 a.
If the total number of bosons is larger than 1 and is even, the system’s ground state is represented by a simple tensor product of the local Fock states with exactly $\frac{N}{2}$ bosons per site Ultracold . There is no degeneracy of the ground state in this case. In turn, if $N$ is odd, the ground state is two-fold degenerate and the entanglement properties are similar to the $N$=1 case.
VII Conclusion
@@
In this work, we address the problem of characterizing quantum systems in degenerate ground states with a special focus on the properties of entanglement entropy. On the basis of the linearly independent degenerate eigenstates calculated with exact diagonalization we propose to generate a finite ensemble of their linear combinations with Haar measure. This allows to probe the entanglement distributions of the ground wave functions. The primers we consider in this work show that such distributions are problem specific. For instance, in the case of the transverse-field Ising model the calculated values of the von Neumann entropy are non-uniformly distributed in the wide range between 0 and 1. At the same time, for the quantum system with non-collinear order the entanglement distribution reveals a more localized Gaussian-like shape. The found enhancement of the mean entropy of the degenerate eigenstate ensemble in the spin spiral phase can be explained by the changes in three- and two-spin correlations which demonstrate different trends at increasing the external magnetic field. To establish a connection between our theoretical results and real experiments we analyze the important details concerning estimation of the observables in the case of the degenerate quantum systems.
Acknowledgments
@@
We thank Stanislav Straupe, Aleksey Fedorov and Evgeniy Kiktenko for fruitful discussions. The research funding from the Ministry of Science and Higher Education of the Russian Federation (Ural Federal University Program of Development within the Priority-2030 Program, project 4.72) is gratefully acknowledged. The quantum VQE simulations were performed within the Russian Roadmap on Quantum Computing (Contract No. 868-1.3-15/15-2021, October 5, 2021) in the development of the trapped-ion processor. The investigations of the fermionic and bosonic models were carried out with the financial support of the Ministry of Science and Higher Education of the Russian Federation (theme FEUZ-2023-0013). Quantum simulations were performed on the Uran supercomputer at the IMM UB RAS.